OpenStudy (anonymous):

simplify (sin x - cos x )^2

OpenStudy (luigi0210):

Distribute: \[\LARGE (sinx-cosx)(sinx-cosx)\]

OpenStudy (anonymous):

sin^2x-sinxcosx-cosc sinx +cos^2x

OpenStudy (isaiah.feynman):

Correct.

OpenStudy (luigi0210):

Alright, now keep going

OpenStudy (anonymous):

sin^2x+cos^2x

OpenStudy (luigi0210):

yes, and what does that become?

OpenStudy (anonymous):

1

OpenStudy (luigi0210):

Now combine the remaining terms

OpenStudy (anonymous):

how

OpenStudy (luigi0210):

Well you have this left: \[\LARGE 1-sinxcosx-sinxcosx\]

OpenStudy (luigi0210):

If it helps, think of the sinxcosx as x's

OpenStudy (luigi0210):

So: 1-x-x

OpenStudy (anonymous):

:)

OpenStudy (luigi0210):

S what is your final annswer?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

When we combine them, we get: 1-2sinxcosx, and 2sinxcos=sin(2x) So final answer: \[\LARGE 1-\sin(2x)\]

OpenStudy (anonymous):

thank you so so much :)

OpenStudy (luigi0210):

Yup!

OpenStudy (luigi0210):

btw: \[\LARGE \mathbb{\color{#00FFCC}{WELCOME~TO~OPENSTUDY}} {\ddot\smile} \]